Operation Summary for September 28, 2011 - September 30, 2011 |
Wednesday September 28 The day shift began with Operations monitoring the Tevatron (TEV) store 9155 (81.18E30) and the Recycler stash (375.5E10), with the Antiproton Source (Pbar) stacking (12.1E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, a Booster person began a study at 8:36 AM. Operations terminated store 9155 at 9:05 AM, put the TEV into HEP shot setup, and established store 9156 at 10:10 AM with an average initial luminosity (AIL) of 347.44E30. At 2:03 PM, operators reported seeing and hearing ICW pipes in the Cross Gallery area shaking violently; FESS personnel de-energized the system. During the evening shift, operators monitored the accelerator complex. Thursday September 29 The midnight shift began with Operations monitoring store 9156 (79.1E30) and the stash (362E10), with Pbar stacking (9.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 12:20 AM, a MI RF station (MIRF13) tripped off and would not reset; operators contacted a RF person who said its modulator needed replacing. Operators bypassed the station. Operations terminated store 9156 at 1:56 AM, put the TEV into shot setup, and established store 9157 at 2:46 AM with an AIL of 373.61E30. During the day shift, Operators rebooted the MI low-level RF system at 8:44 AM; this held off beam for about 35 minutes. A Booster person conducted a study beginning at 10:15 AM. RF personnel repaired MIRF13 and brought it back online. During the evening shift, Operations received the Operational Readiness Clearance at 5:08 PM for the next Meson FTBF experiment T-1012. At 7:12 PM, all the MI abort gap kickers tripped off; they all reset. Operations established beam to T-1012 at 7:48 PM. Operations terminated store 9157 around 7:52 PM, put the TEV into shot setup, and established store 9158 at 9:05 PM with an AIL of 362.4E30. Operations disabled beam to T-1012 at 9:58 PM. Friday September 30 The midnight shift began with Operations monitoring store 9158 (191.58E30) and the stash (115.8E10), with Pbar stacking (20E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators reported at midnight that they had lost communications with the TEV vacuum system from sectors D1 through to F2; operators temporarily bypassed these alarms. Operations halted beam to NuMI at 7:04 AM in preparation for a scheduled access. Operations established beam to T-1012 at 7:38 AM. The Plans for Friday and the weekend The plans for the day are to keep store 9158 in the TEV until 2 PM, which is when the Tevatron will be shut off forever. Beam to NuMI, MiniBooNE, and the Meson FTBF experiments will continue as normal. CDF will go into access. D0 intends to conduct calibration runs and take cosmic rays for a month or so. The Operations Department would like to say so long to the Tevatron. |
|
More Information Comments and Suggestions |
| last modified 09/30/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |